OpenStudy (anonymous):

Match each equation to how it changes the parent function y=x2 (x2 will be written as x^2) this is the problem : y = (x-8)^2

OpenStudy (anonymous):

i think is slides down because of the -2 but im not sure

ganeshie8 (ganeshie8):

thats a very good guess! but its wrong here you're not subtracting anything directly from `y value`, so we cannot say it slides down

OpenStudy (anonymous):

would it possibly flip down instead of sliding?

ganeshie8 (ganeshie8):

think of it like this : y = (x-8)^2 \(\large x = \pm\sqrt{y} \color{red}{+ 8}\)

ganeshie8 (ganeshie8):

you're adding "8" to teh `x value` so it slides 8 units to the right

OpenStudy (anonymous):

so it would go up due to the +8 and not minus

ganeshie8 (ganeshie8):

you're adding 8 to the "x" value, so it moves 8 units forward in "x" direction

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):